OpenStudy (anonymous):

Adina has a rectangular garden that measures 9 m wide by 13 m long. She wants to increase the area to 192 m2 by increasing the width and length by the same amount. What will be the width (shorter dimension) of the new garden?

OpenStudy (anonymous):

@hartnn do you have time for another?

hartnn (hartnn):

let she increase both by 'x' m new width = 9+x so, (9+x)(13+x) = 192 solve for x

hartnn (hartnn):

i hope you got the answer as 12 m

OpenStudy (anonymous):

I did Thanks!!

OpenStudy (anonymous):

That was the right answer right? That's what i got.

hartnn (hartnn):

yes. x=3 new width = 3+9 =12